OpenStudy (anonymous):

If f(2) = 3 and f '(2) = 3, find an equation of the following at the point where x = 2

OpenStudy (anonymous):

they give you the coordinates (2, 3) when they give you f^1(2) = 3, 3 is the slope. so do y=m(x-xp) + yp to find the equation y = 3(x-2) + 3 y = 3x - 6 + 3 y = 3x-3 i hope that's right... haha not 100%